OpenStudy (anonymous):

The zero-product property states that for all real numbers a and b, if a • b = 0, _____.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Zero-product_property A and/or B must equal 0.

OpenStudy (anonymous):

Proof: Suppose that a = 0. Then 0 * b = 0 by the multiplicative identity property. Suppose that b = 0. Then a * 0 = 0 by the multiplicative identity property. Suppose both a and b are 0. Then 0 * 0 = 0 by the multiplicative identity. QED. If this helps, please leave me a medal. http:/www.tutorsean.net

OpenStudy (anonymous):

then a and/or b = 0